Hello.... What is order of magnitude ?.....Can u explain it breifly

Hello....
    What is order of magnitude ?.....Can u explain it breifly

Grade:9

1 Answers

Umakant biswal
7 years ago
@ jeeva 
i am explaining it through an example , try to understand 
suppose u need to find the order of magnitude of 385000 kg 
so, 
u knows , 385000 = (3.85 * 10 ^ 2 ) 10 ^ 3 
= 3.85 * 10 ^ 3+2 
= 3.85 * 10 ^ 5 
hence , 3.85 > 3.2 , so, order of magnitude = 10 ^ 5+1 
order of magnitude = 10 ^ 6 
HOPE IT CLEARS